# Tracing Galaxy Evolution in the Nearby Universe: The Role of Dark Matter

### v1

@@ V.5 Comparison with results from the literature
On the other hand, large uncertainties affected our data do not erase this systematic behavior: this affirmation can be suported following Nigoche-Netro et al. (2009). In appendix A of that paper they shown, using bootstrap resampling and hypothesis testing, that even when uncertainties are comparable to a particular trend amplitude, the monotonic trend remains statistically significant.
Finally, we calculated the logarithmic dynamical-to-baryonic mass difference for the Milky Way and Andromeda (M31) inside their optical radii, obtaining $\Delta$ $\log(\mathbf{M_{MW}})=0.54\pm 0.15$ dex and $\Delta$ $\log(\mathbf{M_{M31}})=0.53\pm 0.15$ dex. The detailed calculation of these parameters is presented in Appendix A (Section A.12). The positions of these two Local Group galaxies in Figures A2 and A3 (blue dot for the Milky Way and red dot for Andromeda) fall within the trends obtained in this work for galaxies of comparable dynamical mass and redshift. The results show that the dark matter inside the optical radius of both galaxies is greater than that observed for galaxies of similar mass at high redshift. This finding reinforces that our methodology and conclusions regarding the mass difference at higher redshifts are consistent with well-characterized nearby galaxies.
@@ VI Conclusions
In this study, we utilize a sample of approximately 126,000 LTGs from the SDSS DR16, spanning the redshift range of approximately $0.00<z<0.35$ and dynamical mass range of $9.5<\log(\mathbf{M_{Dyn}/M_{Solar}})<12.5$. We conduct a comparison between the stellar mass and the dynamical mass, considering eight stellar mass estimates obtained through different population synthesis models and different initial mass functions. Additionally, we consider seven dynamical mass estimates based on Newtonian dynamics, virial equilibrium, different components (gas, stars), and various variables that may impact the estimation of dynamical mass (galactic inclination angle, colour, concentration index, scale factor $K$, Sérsic index). The calibration of dynamical masses was performed following the methodology outlined by Nigoche-Netro et al. (2022). The analysis has yielded the following results:
@@
  • The difference between dynamical and stellar mass, in general, appears to decrease as redshift rises (see Fig. 4 and the figures in Appendix B). This behavior is similar when considering different variables in the dynamical mass (LTGs dominated by the disk or by the bulge) and stellar mass (different population synthesis models, different IMF) estimations. However, this behavior appears to be more pronounced for those LTGs dominated by the disk.
  • The difference between dynamical and stellar mass can range from nearly zero to approximately 95 percent of the dynamical mass, depending on dynamical mass and redshift (refer to Fig. 5).
  • The behavior of the difference between dynamical and stellar mass as a function of dynamical mass and redshift reveals a complex pattern (see Fig. 5 and figures in appendix C). At low dynamical mass and low redshift, the difference is relatively large and tends to decrease as dynamical mass increases, producing a saddle-like shape. This saddle is more pronounced in disk-dominated LTGs, though it is also present in bulge-dominated systems. As redshift increases, this saddle-like behavior gradually transitions into a steeper, more linear trend, where the mass difference increases with dynamical mass. In the high-mass and high-redshift regime, the relation approximates a straight line with positive slope, resembling the trend reported for early-type galaxies (ETGs) by Nigoche-Netro et al. (2016). In this regime, the mass difference grows with dynamical mass, while overall decreasing with redshift.
@@
    These findings appear to indicate that the influence of dark matter in galaxy formation and evolution is more complex than previously understood. They suggest that models of galaxy formation and evolution may need to incorporate a more seamless transition between different morphological types.
@@
We thank the Instituto de Astronomía y Meteorología (UdG, México) for all the facilities provided for the realisation of this project. A. Nigoche-Netro acknowledges support from CONAHCyT and PRODEP (México). Eduardo de la Fuente thanks Colegio departamental de Física, authorities of CUCEI, and Coordinación General Académica y de Innovación (CGAI-UDG). P. Lagos gratefully acknowledges support by the GEMINI ANID project No. 32240002.
Appendix A Variables, biases, and corrections in the estimation of dynamical, stellar and dark matter in galaxies
Estimation of the dynamical and stellar masses of LTGs is a fundamental aspect of our analysis. Below, we provide a comprehensive compilation of the variables, corrections, and potential biases that may affect dynamical, stellar and dark matter estimations:
@@
    Stellar and non-stellar (gas) components. $\sigma$ may be influenced by the presence of stellar and non-stellar components. This bias is linked to the galaxies’ structure, colour, and morphological characteristics.
@@
    Inclination angle. The orientation of the galaxy relative to the observer can affect measurements of effective radius.
A.3 Initial Mass Function (IMF): Determines how stellar masses are distributed at the time of their formation.
— Biases:
@@
A.4 Colour Indices: Used to infer the stellar population and the age of the galaxy.
— Biases:
@@
    Dust extinction affects colour indices, which can lead to errors in the estimation of stellar mass.
@@
    Metallicity influences colour indices, potentially causing inaccuracies in the stellar mass estimation.
@@
    Models using different IMFs can produce variations in stellar mass. This bias is related with A.3(a).
@@
    Models using different star formation history can result in differences in stellar mass.
@@
    Models underestimate the baryonic mass of galaxies because they do not completely account for the weak and non-luminous baryonic matter.
A.7 The factor $K$ (refer to equation 4): This factor indicates whether the proportion between mass and velocity dispersion is constant or varies among different systems or settings. It influences the calculation of the dynamical mass.
— Biases:
@@
    $K$ could depend on the Sérsic index (morphology) (Cappellari et al., 2006; Mocz et al., 2012).
@@ A.9 Other Biases:
    Sample completeness: This factor refers to the extent to which the surveys include all galaxies within its target range of brightness (Malmquist bias), distance, and other selection criteria, ensuring that the sample accurately represents the overall galaxy population. This bias is related to A.5(c).
@@ A.10 Possible dark matter dependencies:
    Dynamical or luminous (stellar) mass: The amount of dynamical or luminous (stellar) mass could be related with the content of dark matter inside galaxies.
@@
    Redshift: The evolution of galaxies over the cosmic time could be related with the content of dark matter inside galaxies.
@@
    Density of galaxies: The environment could be related with the content of dark matter inside galaxies.
The previously listed factors are crucial for obtaining accurate estimates of the dynamical, stellar and dark matter of galaxies. Proper corrections and awareness of these biases are essential for the interpretation of astrophysical data. In the following sections, we will address the correction and biases that can affect mass estimations.
@@ A.11 Correction of Photometric and Spectroscopic Data
  • Seeing and Extinction Corrections: We utilize seeing-corrected parameters for the total magnitude and effective radius, along with extinction corrections, by applying the corresponding SDSS pipelines to the data (see York et al., 2000; Blanton et al., 2003, and references therein). This paragraph considers the biases discussed in Sections A.2(a), A.2(c), A.8(a), and A.8(b).
  • K Correction: The values are derived using the following formulae (Nigoche-Netro et al., 2008):
@@
    This paragraph considers the bias discussed in Section A.5(a).
  • Cosmological Dimming Correction. According to Jørgensen et al. (1995a), the effective surface brightness ($\mu_{e}>$) or magnitude are adjusted by subtracting 10 $\log(1+z)$. This paragraph considers the bias discussed in Section A.8(c).
@@
    This paragraph considers the bias discussed in Section A.8(d).
@@
    This paragraph considers the bias discussed in Section A.2(b).
@@
    where $r_{ap}$ = 1.5 arcsec for SDSS data (York et al., 2000; Blanton et al., 2003).
    This paragraph considers the bias discussed in Section A.1(d).
@@ A.12.2 Mass modeling
The dark matter mass was computed assuming an NFW (Navarro–Frenk–White) density profile (Navarro et al., 1997):
@@
For the Milky Way, we adopted the parameters derived by Petač (2020) and Sofue (2012)
  • Scale density: $\rho_{s}=0.012~M_{\odot}~\mathrm{pc}^{-3}$
  • Scale radius: $r_{s}=18~\mathrm{kpc}$
  • Total baryonic mass: $M_{\text{bar,tot}}=(6.0\pm 1.2)\times 10^{10}~M_{\odot}$
@@
For Andromeda, we adopt dynamical and baryonic masses consistent with the modelling of Tamm et al. (2012). Tamm et al. (2012) report a total stellar mass in the range $(1.0\textrm{--}1.5)\times 10^{11}\,M_{\odot}$ (see their Stellar mass distribution in Table 2 and Discussion). The observed rotation curve and enclosed mass table (see their Table 3 and Fig. 5) imply cumulative masses of order a few $10^{11}\,M_{\odot}$ within radii $\sim$ 20–30 kpc. Adopting
@@
Figure A2: Difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-ED)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{ED}/{\bf M_{Solar}}})$) as function of dynamical mass for the LTGs samples. Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). Each colour and symbol represents a quasi-constant redshift. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{ED}/{\bf M_{Solar}}})$ is approximately 0.243 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (Black circles $z\sim 0.025$), intermediate (purple squares $z\sim 0.095$), and high (dark green triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point. The blue circle represents the Milky Way, and the red circle represents the Andromeda galaxy. The orange bars represent the boundaries of the quasi-constant mass bins shared by the $z=0.095$ and $z=0.165$ redshift samples.
Figure A3: Mosaic of the difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-ED)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{ED}/{\bf M_{Solar}}})$) as a function of dynamical mass, analogous to Fig. 5. Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). In the first column, the quasi-constant redshift trend lines are spaced every 0.02 instead of 0.01, which reduces crowding and makes the overall behavior easier to visualize. In the second column, three representative redshifts are shown: low light green diamonds $z\sim 0.045$), intermediate (magenta triangles $z\sim 0.085$), and high (brown diamonds $z\sim 0.145$). These examples complement those shown in Fig. 5, reinforcing the robustness of the behaviors discussed and facilitating the identification of completeness ranges shared by the different trends. The blue circle represents the Milky Way, and the red circle represents the Andromeda galaxy. The orange bars represent the boundaries of the quasi-constant mass bins shared by the $z=0.085$ and $z=0.145$ redshift samples.
@@
Figure A6: Distribution of dynamical mass ($\log(\mathbf{M_{IB}/M_{Solar}})$) as a function of redshift. Extracted from the last row of Figure 7, this panel visually outlines our binning strategy in dynamical mass and redshift. It illustrates the overlapping mass range for three quasi-constan redshifts: light green for $z=0.045$, magenta for $z=0.085$, and brown for $z=0.145$ (the color scheme is consistent with Figures 6 and A3). Two key regions are indicated: Zone 1: The region dominated by Malmquist bias, where lower-mass galaxies become progressively undetectable at higher redshifts. Zone 2: The region where the survey is sensitive enough to detect galaxies.
@@
Figure A7: Distribution of dynamical mass ($\log(\mathbf{M_{IB}/M_{Solar}})$) as a function of redshift. Extracted from the last row of Figure 7, this panel visually outlines our binning strategy. It highlights a specific quasi-constant dynamical mass bin (centered at $\log(\mathbf{M_{IB}/M_{Solar}})=11.5$, width of $0.1$ dex) across three quasi-constant redshifts: $z=0.045$ (light green), $z=0.085$ (magenta), and $z=0.145$ (brown). The color scheme is consistent with Figures 6 and A3. Two key regions are indicated: Zone 1: The area dominated by Malmquist bias, where lower-mass galaxies become progressively undetectable at higher redshifts. Zone 2: The region where the survey is sensitive enough to detect galaxies.
Appendix B Stellar mass vs redshift

### v2

@@ V.5 Comparison with results from the literature
On the other hand, the large uncertainties affecting our data do not erase this systematic behavior; this affirmation can be supported following Nigoche-Netro et al. (2009). In Appendix A of that paper, they showed, using bootstrap resampling and hypothesis testing, that even when uncertainties are comparable to a particular trend amplitude, the monotonic trend remains statistically significant.
Finally, we calculated the logarithmic dynamical-to-baryonic mass difference for the Milky Way and Andromeda (M31) inside their optical radii, obtaining $\Delta$ $\log(\mathbf{M_{MW}})=0.54\pm 0.15$ dex and $\Delta$ $\log(\mathbf{M_{M31}})=0.53\pm 0.15$ dex. The detailed calculation of these parameters is presented in Appendix A.12. The positions of these two Local Group galaxies in Figures A2 and A3 (blue dot for the Milky Way and red dot for Andromeda) fall within the trends obtained in this work for galaxies of comparable dynamical mass and redshift. The results show that the dark matter inside the optical radius of both galaxies is greater than that observed for galaxies of similar mass at high redshift. This finding reinforces that our methodology and conclusions regarding the mass difference at higher redshifts are consistent with well-characterized nearby galaxies.
@@ VI Conclusions
In this study, we utilize a sample of approximately 126,000 LTGs from the SDSS DR16, spanning the redshift range of approximately $0.00<z<0.35$ and dynamical mass range of $9.5<\log(\mathbf{M_{Dyn}/M_{Solar}})<12.5$. We conduct a comparison between the stellar mass and the dynamical mass, considering eight stellar mass estimates obtained through different population synthesis models and different IMFs. Additionally, we consider seven dynamical mass estimates based on Newtonian dynamics, virial equilibrium, different components (gas, stars), and several variables that may impact the estimation of dynamical mass (galactic inclination angle, color, concentration index, scale factor $K$, Sérsic index). The calibration of dynamical masses was performed following the methodology outlined by Nigoche-Netro et al. (2022). The analysis has yielded the following results:
@@
  • The difference between dynamical and stellar mass, in general, appears to decrease as redshift rises (see Fig. 4 and the figures in Appendix B). This behavior is similar when considering different variables in the dynamical mass (LTGs dominated by the disk or by the bulge) and stellar mass (different population synthesis models, different IMF) estimates. However, this behavior appears to be more pronounced for those LTGs dominated by the disk.
  • The difference between dynamical and stellar mass can range from nearly zero to approximately 95% of the dynamical mass, depending on dynamical mass and redshift (see Fig. 5).
  • The behavior of the difference between dynamical and stellar mass as a function of the dynamical mass and redshift reveals a complex pattern (see Fig. 5 and figures in appendix C). At low dynamical mass and low redshift, the difference is relatively large and tends to decrease as dynamical mass increases, producing a saddlelike shape. This saddle is more pronounced in disk-dominated LTGs, though it is also present in bulge-dominated systems. As redshift increases, this saddlelike behavior gradually transitions into a steeper, more linear trend, where the mass difference increases with dynamical mass. In the high-mass and high-redshift regime, the relation approximates a straight line with positive slope, resembling the trend reported for ETGs by Nigoche-Netro et al. (2016). In this regime, the mass difference grows with dynamical mass, while overall decreasing with redshift.
@@
These findings indicate that the influence of dark matter in galaxy formation and evolution is more complex than previously understood. They suggest that models of galaxy formation and evolution may need to incorporate a more seamless transition between different morphological types.
@@
We thank the Instituto de Astronomía y Meteorología (UdG, México) for all the facilities provided for the realization of this project. A.N.-N. acknowledges support from CONAHCyT and PRODEP (México). E.D.L.F. thanks Colegio departamental de Física, authorities of CUCEI, and Coordinación General Académica y de Innovación (CGAI-UDG). P.L. gratefully acknowledges support by the GEMINI ANID project No. 32240002. The work of R.J.D. is supported by the International Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America.
Appendix A Variables, biases, and corrections in the estimation of dynamical, stellar, and dark matter in galaxies
The estimation of the dynamical and stellar masses of LTGs is a fundamental aspect of our analysis. In this appendix, we provide a comprehensive compilation of the variables, corrections, and potential biases that affecting the estimation of dynamical, stellar, and dark matter. Here, we also include figures and tables that provide additional context for the results discussed in the main body of the article (see Figures A1A7 and Table A1).
@@
    Stellar and nonstellar (gas) components. $\sigma$ may be influenced by the presence of stellar and nonstellar components. This bias is linked to the galaxies’ structure, color, and morphological characteristics.
@@
    Inclination angle. The orientation of the galaxy relative to the observer can affect measurements of the effective radius.
A.3 Initial Mass Function: Determines how stellar masses are distributed at the time of their formation.
Biases:
@@
A.4 color Indices: Used to infer the stellar population and the age of the galaxy.
Biases:
@@
    Dust extinction affects color indices, which can lead to errors in the estimation of stellar mass.
@@
    Metallicity influences color indices, potentially causing inaccuracies in the stellar mass estimation.
@@
    Models using different IMFs can produce variations in stellar mass. This bias is related to  A.3(a).
@@
    Models using different star formation histories can result in differences in stellar mass.
@@
    Models underestimate the baryonic mass of galaxies because they do not completely account for the weak and nonluminous baryonic matter.
A.7 The factor $K$ (refer to equation 4): This factor indicates whether the proportion between mass and velocity dispersion is constant or varies among different systems or settings, and influences the calculation of the dynamical mass.
Biases:
@@
    $K$ could depend on the Sérsic index (morphology; Cappellari et al., 2006; Mocz et al., 2012).
@@ A.9 Other Biases:
    Sample completeness: This factor refers to the extent to which the surveys include all galaxies within their target range of brightness (Malmquist bias), distance, and other selection criteria, ensuring that the sample accurately represents the overall galaxy population. This bias is related to A.5(c).
@@ A.10 Possible dark matter dependencies:
    Dynamical or luminous (stellar) mass: The amount of dynamical or luminous (stellar) mass could be related to the content of dark matter inside galaxies.
@@
    Redshift: The evolution of galaxies over the cosmic time could be related to the content of dark matter inside galaxies.
@@
    Density of galaxies: The environment could be related to the content of dark matter inside galaxies.
The previously listed factors are crucial for obtaining accurate estimates of the dynamical, stellar, and dark matter of galaxies. Proper corrections and awareness of these biases are essential for the interpretation of astrophysical data. In the following sections, we will address the corrections and biases that can affect mass estimations.
@@ A.11 Correction of Photometric and Spectroscopic Data
  • Seeing and Extinction Corrections: We utilize seeing-corrected parameters for the total magnitude and effective radius, along with extinction corrections, by applying the corresponding SDSS pipelines to the data (see York et al., 2000; Blanton et al., 2003, and references therein). This paragraph considers the biases discussed in Appendices A.2(a), A.2(c), A.8(a), and A.8(b).
  • K-correction: The values are derived using the following formulae (Nigoche-Netro et al., 2008):
@@
    This paragraph considers the bias discussed in Appendix A.5(a).
  • Cosmological Dimming Correction. According to Jørgensen et al. (1995a), the effective surface brightness ($\mu_{e}>$) or magnitude are adjusted by subtracting 10 $\log(1+z)$. This paragraph considers the bias discussed in Appendix A.8(c).
@@
    This paragraph considers the bias discussed in Appendix A.8(d).
@@
    This paragraph considers the bias discussed in Appendix A.2(b).
@@
[…]
    This paragraph considers the bias discussed in Appendix A.1(d).
@@ A.12.2 Mass modeling
The dark matter mass was computed assuming an Navarro–Frenk–White (NFW) density profile (Navarro et al., 1997):
@@
For the Milky Way, we adopted the parameters derived by Petač (2020) and Sofue (2012).
  • Scale density: $\rho_{s}=0.012~M_{\odot}~\mathrm{pc}^{-3}$;
  • Scale radius: $r_{s}=18~\mathrm{kpc}$;
  • Total baryonic mass: $M_{\text{bar,tot}}=(6.0\pm 1.2)\times 10^{10}~M_{\odot}$.
@@
For Andromeda, we adopt dynamical and baryonic masses consistent with the modeling of Tamm et al. (2012). Tamm et al. (2012) report a total stellar mass in the range $(1.0\textrm{--}1.5)\times 10^{11}\,M_{\odot}$ (see their Stellar mass distribution in Table 2 and Discussion). The observed rotation curve and enclosed mass table (see their Table 3 and Fig. 5) imply cumulative masses of order a few $10^{11}\,M_{\odot}$ within radii $\sim$ 20–30 kpc. Adopting
@@
Figure A2: Difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-ED)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{ED}/{\bf M_{Solar}}})$) as a function of the dynamical mass for the LTGs samples. Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). Each color and symbol represents a quasi-constant redshift. The difference in redshift between consecutive symbols is approximately 0.01. The mean uncertainty of the difference between $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ and $\log({\bf M_{ED}/{\bf M_{Solar}}})$ is approximately 0.243 dex. The left column of the mosaic contains the full range of redshift, while the right column only contains three specific redshifts: low (black circles, $z\sim 0.025$), intermediate (purple squares, $z\sim 0.095$), and high (dark green, triangles $z\sim 0.165$), the latter with the aim of appreciating more clearly the differences in dynamical and stellar masses due to redshift and also to display the uncertainties for each point. The blue circle represents the Milky Way, and the red circle represents the Andromeda galaxy. The orange bars represent the boundaries of the quasi-constant mass bins shared by the $z=0.095$ and $z=0.165$ redshift samples.
Figure A3: Mosaic of the difference between dynamical and stellar mass ($\Delta{\bf logM_{(Dyn-ED)}}$ = $\log({\bf M_{Dyn}/{\bf M_{Solar}}})$ - $\log({\bf M_{ED}/{\bf M_{Solar}}})$) as a function of the dynamical mass, analogous to Fig. 5. Each row corresponds to a specific estimation of dynamical mass ($\mathbf{M_{KS}}$, $\mathbf{M_{KA}}$, $\mathbf{M_{KB}}$, $\mathbf{M_{nS}}$, $\mathbf{M_{IS}}$, $\mathbf{M_{IA}}$, $\mathbf{M_{IB}}$). In the first column, the quasi-constant redshift trend lines are spaced every 0.02 instead of 0.01, which reduces crowding and makes the overall behavior easier to visualize. In the second column, three representative redshifts are shown: low (light green diamonds, $z\sim 0.045$), intermediate (magenta triangles, $z\sim 0.085$), and high (brown diamonds, $z\sim 0.145$). These examples complement those shown in Fig. 5, reinforcing the robustness of the behaviors discussed and facilitating the identification of completeness ranges shared by the different trends. The blue circle represents the Milky Way, and the red circle represents the Andromeda galaxy. The orange bars represent the boundaries of the quasi-constant mass bins shared by the $z=0.085$ and $z=0.145$ redshift samples.
@@
Figure A6: Distribution of dynamical mass ($\log(\mathbf{M_{IB}/M_{Solar}})$) as a function of the redshift. Extracted from the last row of Figure 7, this panel visually outlines our binning strategy in dynamical mass and redshift. It illustrates the overlapping mass range for three quasi-constant redshifts: light green for $z=0.045$, magenta for $z=0.085$, and brown for $z=0.145$ (the color scheme is consistent with Figures 6 and A3). Two key regions are indicated. Zone 1: the region dominated by Malmquist bias, where lower-mass galaxies become progressively undetectable at higher redshifts. Zone 2: the region where the survey is sensitive enough to detect galaxies.
@@
Figure A7: Distribution of dynamical mass ($\log(\mathbf{M_{IB}/M_{Solar}})$) as a function of the redshift. Extracted from the last row of Figure 7, this panel visually outlines our binning strategy. It highlights a specific quasi-constant dynamical mass bin (centered at $\log(\mathbf{M_{IB}/M_{Solar}})=11.5$, width of $0.1$ dex) across three quasi-constant redshifts: $z=0.045$ (light green), $z=0.085$ (magenta), and $z=0.145$ (brown). The color scheme is consistent with Figures 6 and A3. Two key regions are indicated: Zone 1: The area dominated by Malmquist bias, where lower-mass galaxies become progressively undetectable at higher redshifts. Zone 2: The region where the survey is sensitive enough to detect galaxies.
Appendix B Stellar mass versus redshift
This appendix presents a complete set of mosaics illustrating the behavior of stellar mass as a function of the redshift. These mosaics provide supplementary evidence regarding the physical properties of the galaxies analyzed in this study (see Figures B1B7).